# DiffoRA: Enabling Parameter-Efficient LLM Fine-Tuning via Differential Low-Rank Matrix Adaptation

### v1

@@ Abstract
The Parameter-Efficient Fine-Tuning (PEFT) methods have been extensively researched for large language models in the downstream tasks. Among all the existing approaches, the Low-Rank Adaptation (LoRA) has gained popularity for its streamlined design by incorporating low-rank matrices into existing pre-trained models. Though effective, LoRA allocates every module an identical low-rank matrix, which ignores the varying properties and contributions across different components. Moreover, the existing adaptive LoRA solutions rely highly on intuitive importance scoring indicators to adjust the interior rank of the decomposition matrices. In this paper, we propose a new PEFT scheme called DiffoRA, which is theoretically grounded and enables module-wise adoption of LoRA. At the core of our DiffoRA lies a Differential Adaptation Matrix (DAM) to determine which module is the most suitable and essential for fine-tuning. We explain how the designed matrix impacts the convergence rate and generalization capability of a pre-trained model. Furthermore, we construct the DAM via continuous relaxation and discretization with weight-sharing optimizations. We fully implement our DiffoRA and design comprehensive experiments to evaluate its performance. The experimental results demonstrate that our approach achieves the best model accuracy over all the state-of-the-art baselines across various benchmarks.
@@ 1 Introduction
In recent years, large language models (LLMs) have gained significant attraction across various domains of natural language processing, demonstrating remarkable capabilities in tasks such as text generation [18, 25], machine translation [33], sentiment analysis [32], and question-answering [40]. Due to the large model size, fine-tuning an LLM to some downstream tasks will invoke a large number of parameters, e.g., up to 175 billion parameters for GPT-3 in a fully fine-tuning. As a result, Parameter-Efficient Fine-Tuning (PEFT) becomes increasingly paramount due to its alignment with the LLM demands.
@@
A plethora of PEFT methods have been proposed, which can be categorized into two classes. Some approaches slightly modify the model structure by adding small trainable modules and keeping the rest of the models unchanged [19, 16]. Another line of work aims to efficiently capture the incremental parameter updates without modifying the model structure [7]. Among all the existing work, the Low-Rank Adaptation [12] is widely acknowledged due to its effectiveness and satisfying performance. Unlike the previous work, LoRA incurs some low-rank decomposition matrices to parameterize the incremental updates and realizes comparable results with 70% less overhead than the full fine-tuning. This innovative approach paves the way for more effective utilization of LLMs in practice.
Though effective, there are some limitations of LoRA-based methods. The most fundamental one is that LoRA treats all the modules in the network equally by adding the decomposition matrices to all the trainable modules in the network. Consequently, it ignores the varying properties and contributions of different modules in fine-tuning. Some works have been proposed to adaptively utilize LoRA. Generally speaking, these approaches managed to score the importance of different modules and adjust the interior rank of each decomposition matrix accordingly [37, 36, 21]. For example, Zhang et al. [36] designed AdaLoRA, which utilizes the singular value of the low-rank matrices to adjust the rank. These methods rely highly on the intuitive metrics (e.g., singular values or norms [23, 38]) to score the importance across different modules, which overlook theoretical relationships between the evaluation metrics and the incremental updates of the model. Moreover, the performance of the existing methods can be further enhanced.
@@
Our Objectives. To alleviate the above limitations, in this paper, we aim to answer the following key question: Can we construct a theoretically grounded PEFT method to enable adaptively adoption of the low-rank decomposition matrices, thus we can only fine-tune the modules that are most necessary and essential?
Our contributions. To this end, in this paper, we shed some light on improving the PEFT performance based on LoRA and propose a novel method called DiffoRA. We argue that instead of adjusting the interior rank of every decomposition matrix like previous works, it suffices to adopt the low-rank matrices module-wisely. Thus, at the core of our approach lies a Differentiable Adaptation Matrix (DAM) that determines the importance of each module, so that which one needs to be fine-tuned and vice versa. Unlike the previous intuitive metrics, DAM is processed to be differentiable w.r.t. the incremental updates of the low-rank matrices, thus enabling the capture of the essential characteristics of the module importance. Specifically, we first analyze how the DAM impacts the model performance (i.e., the convergence rate and generalization capabilities) during fine-tuning in theory. We then elaborate on the algorithm to approximate the NP-hard problem by constructing the DAM via continuous relaxation and discretization. To alleviate the possible discrepancy problem and enhance the robustness of DiffoRA, we further incorporate the weight-sharing strategy as optimization. We fully implement our DiffoRA and design extensive experiments to evaluate its performance on two widely adopted benchmarks and multiple tasks. The experimental results show that our scheme is consistently better than all the existing baselines. For instance, DiffoRA achieves 0.81% higher model accuracy on CoLA task [31] than the state-of-the-art method. In all, the contributions of this paper are as follows.
@@
  • We propose DiffoRA, a novel PEFT method for LLMs that is adaptive and theoretically grounded. Our approach is built atop a newly designed differentiable matrix (i.e., DAM), which enables adaptive adoption of the low-rank decomposition matrices.
@@
  • We fully implement our DiffoRA and evaluate its performance on two widely adopted benchmarks that contain multiple tasks. The experimental results demonstrate that DiffoRA works consistently better than all the baselines.
@@ 2 Related work
2.1 LLM and fine-tuning
LLM has captured considerable public interest due to its success in multiple regions. The PEFT is essential for LLMs due to the huge amount of parameters. Some previous works have been proposed to fine-tune the LLMs using specifically designed modules that are added to LLMs. The fine-tuning of LLMs is thus converted to the adjustments of these small modules. For instance, multiple methods [28, 19, 16] insert dataset-dependent small modules or vectors between the layers to decrease the parameter amounts. Another line of work models the incremental updates of the fine-tuning procedure to make it parameter-efficient. Guo et al. [7] propose to use a task-specific difference vector to extend and fine-tune the base model. There are also some methods that fine-tune parts of the parameters [6], e.g., the bias of FFN [35] or the last quarter [15].
@@ 2.2 Low-rank adaptation and optimizations
To further reduce the computational and storage cost, Hu et al. [12] proposed LoRA, in which they designed a low-rank decomposition matrix to model the incremental updates during fine-tuning. A plethora of work has been proposed to optimize LoRA and reduce the parameter amount [14, 8, 34, 2, 29]. One of the limitations of LoRA is that it treats all the modules equally, which omits the variations of the modules in LLMs. To address this issue, a few works have been proposed to realize an adaptive LoRA. Without loss of generality, these methods first evaluate the importance of the modules and then adjust the interior rank of the decomposition matrices accordingly. For instance, AdaLoRA [36] utilizes the singular value to score the importance, while some other approaches adopt the Frobenius norm [23] or norm of the outputs [38] as the metrics. These indicators are intuitively adopted and only utilize partial information of the incremental updates. There are also a few works that utilize the training procedures to determine the module importance [3, 37, 21], yet they still focus on modifying the interior ranks of the decomposition matrices.
@@ 3.1 Preliminaries
Notations. We define $[n]=\{1,2,\ldots,n\}$. We denote the vectors and matrices as the lower and uppercase bold font, respectively. For instance, $\bm{x}$ is a vector with entry $x_{i}$, and $\bm{M}$ is a matrix with entry $[\bm{M}]_{ij}$. The minimum eigenvalue of $\bm{M}$ is denoted as $\lambda_{\min}(\bm{M})$. $\|\cdot\|_{2}$ is used to represent the $l_{2}$ norm of a vector. $N(\bm{0},\bm{I})$ and $U\{S\}$ represent the standard Gaussian distribution and uniform distribution over a set $S$, respectively. We denote by $\bm{X}=\{(\bm{x}_{i},y_{i})|\bm{x}_{i}\in\mathbb{R}^{d\times 1},y_{i}\in\mathbb{R},i\in[n]\}$ the training set, where $\bm{x}_{i}$ and $y_{i}$ represent the $i$-th data and label. $\mathbb{I}\{\cdot\}$ represents the indicator function that demonstrates the event occurrence, such that for event $\mathcal{A}$, $\mathbb{I}\{\mathcal{A}\}=1$ if and only if $\mathcal{A}$ happened, otherwise it equals to 0. $P(\mathcal{A})$ represents the probability of $\mathcal{A}$ occurred event.
Neural networks and gram matrix. For input $\bm{x}\in\mathbb{R}^{d\times 1}$, weight vector $\bm{w}\in\mathbb{R}^{d\times 1}$ in the weight matrix $\bm{W}\in\mathbb{R}^{d\times m}$, and output weight $\bm{a}\in\mathbb{R}^{m\times 1}$, we denote $f(\bm{W},\bm{a},\bm{x})$ as a neural network with a single hidden layer such that
@@
where $\sigma$ is the activation function. In this paper, we primarily consider the ReLU function, which is one of the most adopted activation functions in the literature, i.e., $\sigma(z)=z\mathbb{I}\{z>0\}$. Given a training set $\bm{X}$, the optimization goal is to minimize the empirical risk loss function
@@
In this work, we aim to construct a module-wise DAM $\bm{\Gamma}\in\mathbb{R}^{L\times N}$ with entries $\gamma_{i,j}\in\{0,1\}$ to determine the necessity of each module for fine-tuning in LLM, where $L$ is the number of layers and $N$ is the module amount in each layer. To further analyze the relationship between the model performance and the model with $\bm{\Gamma}$, we expand the definitions in Eq. 1 and Eq. 2 such that
$f(\bm{W},\bm{a},\bm{x};\bm{\Gamma},\bm{W}_{0})=\frac{1}{\sqrt{m}}\sum_{r=1}^{m}a_{r}\sigma((\bm{w}_{0}+\bm{\Gamma}\bm{w}_{r})^{T}\bm{x})$ (3)
$\mathcal{L}(\bm{W},\bm{a};\bm{\Gamma},\bm{W}_{0})=\sum\nolimits_{i=1}^{n}\frac{1}{2}(f(\bm{W},\bm{a},\bm{x}_{i};\bm{\Gamma},\bm{W}_{0})-y_{i})^{2}$
where $\bm{w}_{i}$ is the $i$-th row of $\bm{W}$. Furthermore, we follow the definitions in [4] and define the matrices $\mathbf{H}^{\infty}_{\Gamma,w_{0}}$ and $\mathbf{H}^{\infty}_{w_{0}}$ based on $\bm{\Gamma}$ such that
Definition 1 (Gram Matrix).
For a neural network with a single hidden layer, the gram matrix $\mathbf{H}^{\infty}_{\Gamma,w_{0}}\in\mathbb{R}^{n\times n}$ induced by the ReLU activation function on a training set $\bm{X}:=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ with entry
$[\bm{H}_{\Gamma,w_{0}}^{\infty}]_{ij}=$ $\mathbb{E}_{\bm{w}\sim N(\bm{0},\bm{I})}[\bm{x}_{i}^{T}\bm{x}_{j}\cdot$ (4)
$\mathbb{I}\{(\bm{w}_{0}+\bm{\Gamma}\bm{w})^{T}\bm{x}_{i}\geq 0,(\bm{w}_{0}+\bm{\Gamma}\bm{w})^{T}\bm{x}_{j}\geq 0\}]$
@@
We further construct $\mathbf{H}_{w_{0}}^{\infty}$ with entry $[\mathbf{H}_{w_{0}}^{\infty}]_{ij}$ such that
$[\bm{H}^{\infty}_{w_{0}}]_{ij}=$ $\mathbb{E}_{\bm{w}\sim N(\bm{0},\bm{I})}[\bm{x}_{i}^{T}\bm{x}_{j}\cdot$ (5)
$\mathbb{I}\{(\bm{w}_{0}+\bm{w})^{T}\bm{x}_{i}\geq 0,(\bm{w}_{0}+\bm{w})^{T}\bm{x}_{j}\geq 0\}]$
@@
where $\bm{I}^{\bm{\Gamma w}}$ and $\bm{I}^{\bm{w}}$ are the expectations of the indicator matrices corresponding to vectors $\bm{\Gamma w}$ and $\bm{w}$, respectively. We denote $\lambda_{0}:=\lambda_{\min}(\mathbf{H}_{w_{0}}^{\infty})$, and $\lambda_{0}^{\Gamma}:=\lambda_{\min}(\mathbf{H}_{\Gamma,w_{0}}^{\infty})$.
Recall LoRA. LoRA [12] utilizes two matrices $\bm{A}\in\mathbb{R}^{r\times k},\bm{B}\in\mathbb{R}^{d\times r}$ to substitute the parameters’ increments. For $\bm{h}=\bm{W}^{(0)}\bm{x}$, the forward pass in LoRA is
$\bm{h}=\bm{W}^{0}\bm{x}+\Delta\bm{x}=\bm{W}^{0}\bm{x}+\bm{B}\cdot\bm{A}\bm{x}$ (6)
where $d,k\ll r$, $\bm{A}$ is usually initialized by following Gaussian distribution and $\bm{B}$ is initialized with zeros. LoRA adopts this modification equally to all the modules in the model. In the following, we denote $\Delta\bm{W}$ as $\bm{B}\cdot\bm{A}$.
3.2 Main theorems
Technical intuitions. In this section, we focus on analyzing the model performance regarding two aspects, i.e., the convergence rate and the generalization capability. The existing schemes that adaptively adopt LoRA focus on adjusting the interior rank of the matrices based on some intuitive evaluation metrics. In contrast to these methods, we argue that adopting LoRA module-wisely in each layer is sufficient. Thus, our intuition in this work is to exert a binary matrix (i.e., $\bm{\Gamma}$) on the model structure to realize a “selective” fine-tuning of the model. In this section, we first assume there exists an appropriate algorithm to construct $\bm{\Gamma}$ and explain why adopting it can lead to better model performance.
Our analysis is established on the intuitive observations that fine-tuning a pre-trained model can be viewed as training a well-initialized model. We construct the theories to explain how the $\bm{\Gamma}$ of a well-initialized over-parameterized Neural Network (NN) impacts the convergence rate and generalization capability. In the following section, we will first present the theorem regarding $\lambda_{0}$ and the model performance then demonstrate our main theorems.
Theoretical results. The over-parameterized NNs are widely adopted and competitive in hierarchical feature extraction due to the large number of parameters they contain. We use the architecture with wide hidden layers in this section to establish our theories of the matrix $\mathbf{\Gamma}$. This network is one of the most fundamental structures of the over-parameterized NN and is proved to be tractable in training [17, 13]. Based on this insight, we present the following theorem [4] about the convergence rate of the NN with a single hidden layer as follows.

### v2

# DiffoRA: Enabling Parameter-Efficient Fine-Tuning via Differential Module Selection

@@ Abstract
The Parameter-Efficient Fine-Tuning (PEFT) methods have been extensively researched for large language models in downstream tasks. Among all the existing approaches, the Low-Rank Adaptation (LoRA) has gained popularity for its streamlined design by incorporating low-rank matrices into existing pre-trained models. Though effective, LoRA, as well as its adaptive optimizations, either allocate the same matrix to all the modules or adjust the interior rank of the components based on importance scoring indicators. In this paper, we argue that not all the modules in LLMs are suitable and necessary to be fine-tuned. Enlightened by this insight, we propose a new PEFT scheme called DiffoRA, which enables adaptive adoption of the low-rank decomposition matrices. At the core of DiffoRA lies a Differential Adaptation Matrix (DAM) to determine which module is the most suitable and essential for fine-tuning. We theoretically explain how the designed matrix impacts the convergence rate and generalization capability of a pre-trained model. We then construct the DAM via continuous relaxation and discretization with weight-sharing optimizations. We fully implement DiffoRA and design comprehensive experiments to evaluate its performance. The experimental results demonstrate that DiffoRA delivers state-of-the-art results across multiple benchmarks.
@@ 1 Introduction
Large language models (LLMs) have gained significant attention across various domains of natural language processing, demonstrating remarkable capabilities in tasks such as text generation [1, 2], machine translation [3], sentiment analysis [4], and question-answering [5]. Due to the large model size, fine-tuning an LLM to downstream tasks will invoke a large number of parameters, e.g., up to 175 billion parameters to fully fine-tune GPT-3. As a result, Parameter-Efficient Fine-Tuning (PEFT) has become increasingly paramount.
@@
The existing PEFT methods can be categorized into two classes. Some approaches slightly modify the model structure by adding small trainable modules and keeping the rest of the model unchanged [6, 7]. Another line of work aims to efficiently capture the incremental parameter updates without modifying the model structure [8]. Among all the existing work, the Low-Rank Adaptation [9] is widely acknowledged due to its effectiveness and satisfying performance. Unlike the previous work, LoRA incurs some low-rank decomposition matrices to parameterize the incremental updates and realizes comparable results with 70% less overhead than the full fine-tuning. This innovative approach paves the way for more effective utilization of LLMs in practice.
Although LoRA is effective in practice, it treats all the modules in the network equally by adding the same decomposition matrices to all the trainable modules in the network. To alleviate this issue, some work has been proposed as adaptive LoRA, which managed to adjust the interior ranks of decomposition matrices for each module based on parameter importance estimation [10, 11, 12, 13, 14]. For example, Zhang et al. [11] utilizes the singular value of the low-rank matrices to adjust the rank, while SalientLoRA [15] uses a salience measurement to adaptively optimize the ranks of LoRA.
Though effective, we point out that fine-tuning all modules, like the previous work, is not necessary and might inversely affect the model performance. As justification, we conduct a toy experiment by comparing the accuracy of DeBERTaV3-Base on CoLA [16] using LoRA when selectively fine-tuning: S1) all modules; S2) only the QKV projections; S3) only dense modules; S4) QKV projections in the first half of the model and dense modules in the remaining half; and S5) three modules per layer that are randomly selected. We repeat the selection for 10 times to find the optimal configuration. It can be seen from the results in Figure 1 that
  • Different modules and combination strategies contribute differently to the fine-tuning results. An appropriate module combination can significantly enhance the model’s performance. For instance, by randomly selecting the modules in each layer, S5 achieves over 4% higher accuracy than other settings with lower parameter amounts.
  • Not all of the modules are suitable and necessary to be fine-tuned. For instance, S1 obtains a lower accuracy than S4 and S5, which also leads to overfitting. Some modules in S1 contribute negatively to the performance, thus, it is unnecessary to adjust their interior ranks.
@@
The above observations demonstrate the importance of selecting the most essential modules during LoRA fine-tuning, which is often overlooked by the existing methods. This presents a compelling research question: Can we construct a theoretically grounded PEFT method to enable adaptive adoption of the low-rank decomposition matrices, so that we can only fine-tune the modules that are most necessary and essential?
Figure 1: The loss (left) and final accuracy (right) obtained on the CoLA dataset and DeBERTaV3-base using LoRA under different settings.
Our contributions. In this paper, we affirmatively answer the question with DiffoRA, a novel method to improve PEFT performance based on LoRA. Instead of adjusting the interior rank of every decomposition matrix, DiffoRA only adopts the low-rank matrices for those modules that are suitable and necessary to be fine-tuned. To this end, at the core of our approach lies a Differentiable Adaptation Matrix (DAM) that performs a module selection. DAM is designed to be differentiable with respect to the incremental updates of the low-rank matrices, allowing it to effectively capture fine-tuning priorities for various network backbones. We provide a theoretical analysis of how DAM influences convergence rate and generalization during fine-tuning. Then, we approximate the NP-hard optimization by constructing DAM via continuous relaxation and discretization. To mitigate discretization-induced discrepancies and enhance robustness, we further introduce a weight-sharing strategy. We implement DiffoRA and conduct extensive experiments on three standard benchmarks for both Natural Language Understanding and Question Answering tasks. The results demonstrate that DiffoRA consistently outperforms existing methods. For example, it achieves a 1.51% accuracy gain over the state-of-the-art on the CoLA dataset. The main contributions of this work are summarized as follows:
@@
  • We propose DiffoRA, a novel and theoretically-grounded PEFT method for LLMs. Our approach is built atop a newly designed differentiable matrix (i.e., DAM), which enables module-wise adoption of the low-rank decomposition matrices.
@@
  • We fully implement DiffoRA and evaluate its performance on several widely adopted benchmarks that contain multiple tasks. The experimental results demonstrate that DiffoRA works consistently better than all the baselines.
@@ 2 Related work
2.1 Parameter-efficient fine-tuning.
LLMs have captured considerable public interest due to their success in multiple regions. The PEFT is essential for LLMs due to the huge number of parameters. Some previous work has been proposed to fine-tune the LLMs using specifically designed modules that are added to the LLMs. The fine-tuning of LLMs is thus converted to the adjustments of these small modules. For instance, multiple methods [17, 6, 7, 18] insert dataset-dependent small modules or vectors between the layers to decrease the parameter amounts. Another line of work models the incremental updates of the fine-tuning procedure to make it parameter-efficient. Guo et al. [8] propose to use a task-specific difference vector to extend and fine-tune the base model. There are also some methods that fine-tune parts of the parameters [19], e.g., the bias of FFN [20] or the last quarter [21].
@@ 2.2 Low-rank adaptation and optimizations
To further reduce the computational and storage cost, Hu et al. [9] proposed LoRA, in which they designed a low-rank decomposition matrix to model the incremental updates during fine-tuning. A plethora of work has been proposed to optimize LoRA [22, 23, 24, 25, 26, 27, 28, 29]. One of the limitations of LoRA is that it treats all the modules equally, which omits the variations of the modules in LLMs. To address this issue, a few work has been proposed to realize an adaptive LoRA. These methods first evaluate the importance of the modules and then adjust the interior rank of the decomposition matrices accordingly, e.g., AdaLoRA [11] and SalientLoRA [15]. There is also a few work that utilize the training procedures to determine the module importance [30, 10, 12], yet they still focus on modifying the interior ranks of the decomposition matrices.
@@ 3.1 Preliminaries
Notations. We define $[n]=\{1,2,\ldots,n\}$. Vectors and matrices are represented by bold lowercase and bold uppercase fonts, respectively. For instance, $\bm{x}$ is a vector with entry $x_{i}$, and $\bm{M}$ is a matrix with entry $[\bm{M}]_{ij}$. The minimum eigenvalue of $\bm{M}$ is denoted as $\lambda_{\min}(\bm{M})$. $\|\cdot\|_{2}$ is used to represent the $l_{2}$ norm of a vector. $N(\bm{0},\bm{I})$ and $U\{S\}$ represent the standard Gaussian distribution and uniform distribution over a set $S$, respectively. We denote by $\bm{X}=\{(\bm{x}_{i},y_{i})|\bm{x}_{i}\in\mathbb{R}^{d\times 1},y_{i}\in\mathbb{R},i\in[n]\}$ the training set, where $\bm{x}_{i}$ and $y_{i}$ represent the $i$-th data and label. $\mathbb{I}\{\cdot\}$ represents the indicator function that demonstrates the event occurrence, such that for event $\mathcal{A}$, $\mathbb{I}\{\mathcal{A}\}=1$ if and only if $\mathcal{A}$ happened, otherwise it equals 0. $P(\mathcal{A})$ represents the probability of $\mathcal{A}$ occurred event.
For input $\bm{x}\in\mathbb{R}^{d\times 1}$, weight vector $\bm{w}\in\mathbb{R}^{d\times 1}$ in the weight matrix $\bm{W}\in\mathbb{R}^{d\times m}$, and output weight $\bm{a}\in\mathbb{R}^{m\times 1}$, we denote $f(\bm{W},\bm{a},\bm{x})$ as a neural network with a single hidden layer such that
@@
where $\sigma$ is the activation function. In this paper, we primarily consider the ReLU function, which is one of the most widely adopted activation functions in the literature, i.e., $\sigma(z)=z\mathbb{I}\{z>0\}$. Given a training set $\bm{X}$, the optimization goal is to minimize the empirical risk loss function
@@
Recall LoRA. Hu et al. [9] utilize two matrices $\bm{A}\in\mathbb{R}^{r\times k},\bm{B}\in\mathbb{R}^{d\times r}$ to substitute the parameters’ increments, where $d,k\ll r$, $\bm{A}$ is usually initialized by following Gaussian distribution and $\bm{B}$ is initialized with zeros. For $\bm{h}=\bm{W}^{0}\bm{x}$, the forward pass in LoRA is
$\bm{h}=\bm{W}^{0}\bm{x}+\Delta\bm{x}=\bm{W}^{0}\bm{x}+\bm{B}\cdot\bm{A}\bm{x}$ (3)
where $\bm{W}^{0}$ denotes the weight matrix of the pre-trained network, consisting of $r$ row vectors $\bm{w}_{0,r}$. In the following, we denote $\Delta\bm{W}$ as $\bm{B}\cdot\bm{A}$. LoRA adopts this modification equally to all the modules in the model. Therefore, fine-tuning the pre-trained network using LoRA can be formulated as
$f(\bm{W},\bm{a},\bm{x};\bm{W}_{0})=\frac{1}{\sqrt{m}}\sum\nolimits_{r=1}^{m}a_{r}\sigma((\bm{w}_{0,r}+\bm{w}_{r})^{T}\bm{x})$ (4)
where $\bm{w}_{r}$ is the $r$-th row of the trainable parameters in the low-rank matrices.
3.2 Main theorems
Technical intuitions. In this work, we aim to construct a matrix called DAM to determine the necessity of each module for fine-tuning in LLM. We denote DAM as $\bm{\Gamma}$. Intuitively, $\mathbf{\bm{\Gamma}}$ is applied to the trainable parameters of the LoRA module-wise, such that the module corresponding to $\gamma_{i,j}=1$ is determined to be fine-tuned, and vice versa. Notably, the module-wise fine-tuning can be viewed as a special case of an element-wise formulation, where entire submatrices are either enabled or disabled uniformly. For analytical clarity, we focus on the more general element-wise setting, which allows finer-grained control and provides deeper insights into how the structure of $\bm{\Gamma}$ influences models’ performance. In what follows, we assume access to a suitable algorithm for constructing $\bm{\Gamma}$ and aim to elucidate how such selective adaptation contributes to improved model performance. To facilitate this analysis, we extend the definitions introduced in Equation 1 as follows:
$f(\bm{W},\bm{a},\bm{x};\bm{\Gamma},\bm{W}_{0})=\frac{1}{\sqrt{m}}\sum_{r=1}^{m}a_{r}\sigma((\bm{w}_{0,r}+\bm{\gamma}_{r}\circ\bm{w}_{r})^{T}\bm{x})$ (5)
where $\bm{\gamma}_{r}$ is the $r$-th row of $\bm{\Gamma}$ and $\circ$ represents the Hadamard product. Furthermore, we follow the definitions in [31] and define $\mathbf{H}^{\infty}_{\bm{\Gamma},\bm{w}_{0}}$ and $\mathbf{H}^{\infty}_{\bm{w}_{0}}$ based on $\bm{\Gamma}$ such that
Definition 1 (Gram Matrix with $\bm{\Gamma}$).
For a neural network with a single hidden layer and ReLU activation, the Gram matrix $\bm{H}^{\infty}_{\bm{\Gamma},\bm{w}_{0}}\in\mathbb{R}^{n\times n}$ induced on a training set $\bm{X}:=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ is defined entry-wise as
$[\bm{H}_{\bm{\Gamma},\bm{w}_{0}}^{\infty}]_{ij}=$ $\mathbb{E}_{\bm{w}\sim N(\bm{0},\bm{I}),r\in[m]}[\bm{x}_{i}^{T}\bm{x}_{j}\cdot\mathbb{I}\{(\bm{w}_{0,r}+\bm{\gamma}_{r}\circ\bm{w})^{T}\bm{x}_{i}\geq 0,(\bm{w}_{0,r}+\bm{\gamma}_{r}\circ\bm{w})^{T}\bm{x}_{j}\geq 0\}]$ (6)
@@ 3.2 Main theorems
We also construct $\mathbf{H}_{\bm{w}_{0}}^{\infty}$ with entry $[\mathbf{H}_{\bm{w}_{0}}^{\infty}]_{ij}$ such that
$[\bm{H}^{\infty}_{\bm{w}_{0}}]_{ij}=$ $\mathbb{E}_{\bm{w}\sim N(\bm{0},\bm{I}),r\in[m]}[\bm{x}_{i}^{T}\bm{x}_{j}\cdot\mathbb{I}\{(\bm{w}_{0,r}+\bm{w})^{T}\bm{x}_{i}\geq 0,(\bm{w}_{0,r}+\bm{w})^{T}\bm{x}_{j}\geq 0\}]$ (7)
@@
where $\bm{I}^{\bm{\bm{\Gamma}w}}$ and $\bm{I}^{\bm{w}}$ are the expectations of the indicator matrices. We denote $\lambda_{0}:=\lambda_{\min}(\mathbf{H}_{\bm{w}_{0}}^{\infty})$, and $\lambda_{0}^{\bm{\Gamma}}:=\lambda_{\min}(\mathbf{H}_{\bm{\Gamma},\bm{w}_{0}}^{\infty})$.
Theoretical results. The theoretical analysis is based on the intuitive observation that fine-tuning a pre-trained model can be viewed as training a well-initialized model. We analyze LoRA fine-tuning from the perspective of over-parameterized neural networks, focusing on two fundamental aspects: the convergence rate and the generalization capability. Over-parameterized networks have been widely adopted and demonstrated strong empirical performance in hierarchical feature extraction, primarily due to their large capacity and training tractability [32, 33]. In this section, we leverage this architecture to establish our theoretical understanding of the matrix $\mathbf{\bm{\Gamma}}$. Building on prior work [31], we present the following theorem that characterizes the convergence behavior of a single-hidden-layer neural network in the over-parameterized setting.
We first analyze the relationship of the minimum eigenvalue between the adaptive matrix (i.e., $\lambda^{\bm{\Gamma}}_{0}$) and the original matrix (i.e., $\lambda_{0}$). We summarize our results in the following theorem: